# Watermarking Algorithm for Large Language Models

### v1

@@ Abstract
The rapid proliferation of large language models (LLMs) has created an urgent need for reliable methods to detect whether a text is generated by such models. In this paper, we propose SimMark, a posthoc watermarking algorithm that makes LLMs’ outputs traceable without requiring access to the model’s internal logits, enabling compatibility with a wide range of LLMs, including API-only models. By leveraging the similarity of semantic sentence embeddings and rejection sampling to impose detectable statistical patterns imperceptible to humans, and employing a soft counting mechanism, SimMark achieves robustness against paraphrasing attacks. Experimental results demonstrate that SimMark sets a new benchmark for robust watermarking of LLM-generated content, surpassing prior sentence-level watermarking techniques in robustness, sampling efficiency, and applicability across diverse domains, all while preserving the text quality.***The source code of our algorithm can be found at https://github.com/DabiriAghdam/SimMark.
Large Language Models,Watermarking LLMs,Posthoc Watermarking,Sentence-Level Watermarking,Similarity-Based Watermarking, Machine Learning, ICML
@@ 1 Introduction
Figure 1: A high-level overview of SimMark detection algorithm. The input text is divided into individual sentences $X_{1}$ to $X_{N}$, which are embedded using a semantic embedding model, with optional PCA for dimensionality reduction. The similarity between consecutive sentence embeddings is computed. Sentences with similarities within a predefined interval $[a,b]$ are considered valid, while those outside are invalid. A soft-$z$-test is performed using the soft count of invalid/partially valid sentences to determine whether the text is watermarked.
The advent of deep generative AIs has made it increasingly important to detect whether a given text is generated by a large language model (LLM). Advanced LLMs, such as GPT-3.5 (OpenAI, 2022), GPT-4 (OpenAI et al., 2023), and Claude 3 (Anthropic, 2024), can produce human-like text at scale and low cost, enabling transformative applications across industries. However, these capabilities also present significant risks, including but not limited to misuse for academic plagiarism, disinformation campaigns and propaganda aimed at manipulating public opinion. For example, the use of AI-generated content in news articles has raised concerns about transparency, accountability, and the potential spread of misinformation (Futurism, 2023). Furthermore, the ability to reliably identify LLM-generated content is essential for enforcing copyright protections and ensuring accountability for such content (Weidinger et al., 2021).
@@
One promising approach to addressing this, is the use of imperceptible statistical signatures, or watermarks, embedded within a text. Watermarking imperceptibly alters text such that it remains natural to human readers but enables subsequent detection of its origin (Atallah et al., 2001). Effective watermarking methods must balance the preservation of LLM output quality with robustness against adversarial paraphrasing, where attackers modify the text to evade detection (Krishna et al., 2024). Additionally, watermarks must be resistant to spoofing attacks, wherein adversaries craft non-machine-generated text (often malicious) to falsely trigger watermark detectors (Sadasivan et al., 2023).
In this paper, we introduce SimMark, a robust posthoc sentence-level watermarking algorithm for LLMs based on sentence embedding similarity. The term posthoc here refers to the fact that the watermarking process operates on text after it has been generated by the LLM, rather than modifying the text generation process itself. This approach makes SimMark compatible with a wide range of models, including open-weight LLMs and closed-source proprietary models accessible only via APIs, as it does not require fine-tuning or access to the model’s internal logits. Access to logits is often restricted by API providers due to their potential use in distilling LLMs and leaking proprietary information (Finlayson et al., 2024). SimMark leverages embedding representations from text embedding models to capture semantic relationships between sentences and imposes detectable statistical patterns on sentence similarity through rejection sampling. Specifically, rejection sampling involves querying the LLM multiple times until the similarity between embeddings of consecutive sentences falls within a certain predefined interval. At detection time, these patterns are analyzed using one-proportion soft-$z$-test, a statistical test, to differentiate between human-written and LLM-generated text, as illustrated in Figure 1.
@@
  • We introduce a novel watermarking algorithm that achieves state-of-the-art detection performance while maintaining low false positive rates for human-written text.
@@
  • Compared to existing watermarking methods, SimMark provides a more practical solution that operates without access to LLM logits or fine-tuning, offering efficient and high-quality watermark injection and detection.
The remainder of this paper is organized as follows. Section 2 reviews the background and related work on watermarking techniques for LLMs. Section 3 outlines our methodology. Section 4 describes our experimental setup and presents comparative results, while Section 5 concludes the paper. In the end, we also discuss the potential ethical considerations and societal impacts.
@@ 2.1 Autoregressive Decoding of LLMs
An LLM operates over a vocabulary $V$, a set of words or subwords termed as tokens. Let $f:V\rightarrow V$ be an LLM that takes a sequence of tokens $T_{i}=\{t_{1},t_{2},\ldots,t_{i}\}$ as input and generates the next token $t_{i+1}$ as its output.
To generate $t_{i+1}$, the LLM samples it from the conditional probability distribution $P(t_{i+1}|T_{i})$ over the vocabulary $V$. After generating $t_{i+1}$, the updated sequence $T_{i+1}=\{t_{1},t_{2},\ldots,t_{i},t_{i+1}\}$ is fed back into the model $f$, and the process is repeated iteratively to generate the subsequent token $t_{i+2}$. This continues until an end-of-sequence token is generated or the generation reaches a predefined token limit. This sequential process of generating one token at a time based on the previously generated tokens is known as autoregressive decoding.
@@ 2.2 Token-Level Watermarking
Token-level watermarking methods embed a statistical signal in the text by manipulating the token generation process (Kirchenbauer et al., 2023; Zhao et al., 2023; Fu et al., 2024). These methods typically alter the probability distribution over $V$ during decoding, subtly biasing the selection of certain tokens to form detectable patterns.
For instance, the KGW algorithm introduced by Kirchenbauer, Geiping, Wen, Katz, Miers, and Goldstein (2023), groups $V$ into green and red subsets pseudo-randomly seeded on the previous token before generating each new token. A predefined positive constant $\delta$ is added to the logits of each token in the green list, increasing their probability of being selected during the sampling step.
At detection, a one-proportion $z$-test is applied to the number of tokens from the green list in the text to determine whether the text contains a watermark. This test compares the observed proportion of green tokens to the expected proportion under the null hypothesis of no watermark, providing a statistical measure to detect even subtle biases introduced by the watermark (Kirchenbauer et al., 2023).
Zhao et al. (2023)’s UNIGRAM-WATERMARK builds upon KGW by fixing the red and green lists instead of pseudo-randomly selecting them, proving that compared to KGW their method is more robust to paraphrasing and editing (Zhao et al., 2023). However, as outlined by Hou et al. (2024), this algorithm can be reverse-engineered, rendering it impractical for high-stakes, real-world applications.
The Semantic Invariant Robust (SIR) watermark in Liu et al. (2023) is also similar to KGW but is designed to be less sensitive to attacks involving synonym replacement or advanced paraphrasing. SIR achieves this by altering the LLM logits based on the semantics of previously generated tokens, using an embedding model to compute semantic representations and training a model that adjusts LLM’s logits based on the semantic embeddings of prior tokens (Liu et al., 2023).
Detection of such watermarks involves analyzing token sequences for statistical signatures that deviate from typical human-generated text. However, token-level watermarks can still be vulnerable to paraphrasing attacks, as rephrasing may disrupt the green and red token lists without altering the overall meaning (Krishna et al., 2024). Moreover, since these methods modify the logits, they directly impact the conditional probability distribution over $V$, potentially affecting the quality of the generated text (Fu et al., 2024).
@@ 2.3 Sentence-Level Watermarking
One approach to mitigate the previously mentioned problems is to inject the watermark signal at the sentence-level semantics, making it less vulnerable to paraphrasing attacks. Consider a similar notation for sentence generation using model $f$ (i.e., the LLM) with autoregressive decoding that takes a sequence of sentences $M_{i}=\{X_{1},X_{2},\ldots,X_{i}\}$ and generates the next sentence $X_{i+1}$. The updated sequence of sentences $M_{i+1}=M_{i}\cup\{X_{i+1}\}$ is then used to generate subsequent sentences iteratively.
@@
SemStamp (Hou et al., 2023) employs Locality-Sensitive Hashing (LSH) (Indyk & Motwani, 1998) to pseudo-randomly partition the semantic space of the embedding model into a set of valid and blocked regions, analogous to the green and red subsets in KGW. During rejection sampling, if the representation of a newly generated sentence lies within the valid regions (determined based on the LSH signature of the previous sentence), the sentence is accepted. Otherwise, a new sentence is generated until a valid sentence is produced or the retry limit is reached. Similar to KGW, a one-proportion $z$-test is applied to the number of valid sentences (analogous to green tokens) to determine whether the text contains a watermark.
@@
To improve robustness against paraphrasing, SemStamp used a contrastive learning approach (Hadsell et al., 2006), fine-tuning a semantic embedding model such that the embeddings of paraphrased sentences remain as close as possible to the original sentences. This was achieved by minimizing the distance between paraphrased and original embeddings while ensuring unrelated sentences remained distinct. Additionally, they introduce a margin constraint in the rejection sampling process to reject sentences whose representations lie near the region boundaries. Their results demonstrate that this sentence-level approach improves robustness against paraphrasing.
@@
$k$-SemStamp (Hou et al., 2024) builds upon SemStamp and aims to enhance robustness by partitioning the semantic space using $k$-means clustering (Lloyd, 1982) instead of random partitioning. They claim that in this way, sentences with similar semantics are more likely to fall within the same partition, unlike random partitioning, which may place semantically similar sentences into different partitions, reducing robustness (Hou et al., 2024). However, $k$-SemStamp assumes that the LLM generates text within a specific domain to apply $k$-means clustering effectively (Hou et al., 2024), limiting its applicability in real-world, open-domain scenarios. The generation and detection procedures of $k$-SemStamp remain largely similar to the original SemStamp. In contrast to token-level algorithms, these sentence-level methods do not alter the logits of the LLM, therefore it is expected for their resulting text to be of higher quality (Hou et al., 2023).
@@
Our work, like SemStamp and $k$-SemStamp, is a sentence-level watermarking algorithm; however, it embeds its watermark signature in the semantic similarity of consecutive sentences. It achieves greater generalizability across domains by leveraging an off-the-shelf, general-purpose sentence embedding model without fine-tuning. At the same time, it outperforms state-of-the-art token-level and sentence-level watermarking methods in robustness against paraphrasing while preserving text quality.
3 SimMark: A Similarity-Based Watermark
@@
Figure 2: Overview of SimMark algorithm. Top: Generation. For each newly generated sentence ($X_{i+1}$), its embedding ($e_{i+1}$) is computed using the Instructor-Large model, optionally applying PCA for dimensionality reduction. The cosine similarity (or Euclidean distance) between $e_{i+1}$ and the embedding of the previous sentence ($e_{i}$), denoted as $s_{i+1}$, is calculated. If $s_{i+1}$ lies within the predefined interval $[a,b]$, the sentence is marked valid and accepted. Otherwise, rejection sampling generates a new candidate sentence until validity is achieved or the iteration limit is reached. Once a sentence is accepted, the process repeats for subsequent sentences. Bottom: Detection (+ Paraphrase attack). Paraphrased versions of watermarked sentences are generated ($Y_{i}$), and their embeddings ($e^{\prime}_{i}$) are computed. The similarity (or distance) between consecutive sentences in the paraphrased text is evaluated. If paraphrasing causes the similarity ($s^{\prime}_{i+1}$) to fall outside $[a,b]$, it is mismarked as invalid. A soft counting mechanism (via function $c(s_{i+1})$ instead of a regular counting with a step function in the interval $[a,b]$) quantifies partial validity based on proximity to the interval bounds, enabling detection of watermarked text via the soft-$z$-test even under paraphrase attacks. It should be emphasized that soft counting is always applied during detection, regardless of whether paraphrasing is present or not, as we cannot assume prior knowledge of paraphrasing.
@@ 3.1 Watermarked Text Generation
Similar to SemStamp and $k$-SemStamp (Hou et al., 2023, 2024), SimMark utilizes the embedding representations of the sentences. To compute the embeddings, in contrast with Hou et al. (2023, 2024) that fine-tuned their embedder model (which could make it biased toward a specific paraphrasing model or domain), we employ Instructor-Large (Su et al., 2023), a general-purpose instruction-tuned embedding model. The use of a general-purpose embedding model enables our approach to be more easily adaptable to different domains, and since the generated embeddings are conditioned on a specific instruction, even if a third party knows what embedder model we used, it is still not trivial to reverse-engineer our algorithm without knowing the exact instruction used during generation (the user can also change this instruction frequently), resulting in more robustness against spoofing attacks.
Algorithm 1 Generation Pseudo-Code for SimMark
0:  LLM, embedding model, predefined interval $[a,b]$, maximum iterations $N_{\mathrm{max}}$, optional PCA model
1:  for each generated sentence $i$ ($X_{i}$do
@@
5:         Generate sentence $i+1$ ($X_{i+1}$) using the LLM.
@@
8:         Compute similarity $s_{i+1}\leftarrow sim(e_{i},e_{i+1})$:
$\small s_{i+1}\leftarrow\begin{cases}\tfrac{e_{i}\cdot e_{i+1}}{\|e_{i}\|_{2}\cdot\|e_{i+1}\|_{2}}&\text{ if cosine similarity},\\ \|e_{i}-e_{i+1}\|_{2}&\text{ if Euclidean distance,}\end{cases}$
@@
10:     while $s_{i+1}{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}\notin[a,b]}$ and $n<N_{\mathrm{max}}$
11:     Accept $X_{i+1}$ as valid and continue generating the next sentence. {Here we either reached the $N_{\mathrm{max}}$ or $s_{i+1}{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\in[a,b]}$}
@@
Using Instructor-Large, we compute embeddings for a sentence by passing both the sentence and an instruction (e.g., “Represent the sentence for cosine similarity:” or “Represent the sentence for Euclidean distance:”) to the Instructor model. Then, we calculate the cosine similarity (or Euclidean distance) between the embedding of sentence $i+1$ (denoted as $e_{i+1}$ in Algorithm 1) and the embedding of sentence $i$ (denoted as $e_{i}$). If the computed value lies within the user-defined predefined interval, sentence $i+1$ is considered valid (analogous to the green subset in KGW). Otherwise, we prompt the LLM to generate a new sentence and repeat this procedure until a valid sentence is found or the maximum number of iterations is reached (in this case we accept the last generated sentence, and continue to generating sentence $i+2$), as shown in Algorithm 1. Optionally, we may apply Principal Component Analysis (PCA) method to the embeddings to reduce their dimensionality before calculating the similarity or distance (this is the instruction in this case: “Represent the sentence for PCA:”). The reason for applying PCA is provided in Subsection 3.2. An overview of SimMark generation algorithm is depicted in the top part of Figure 2.
The predefined interval is a hyperparameter chosen a priori by the user based on the distribution of similarities (or distances) between consecutive sentences’ embeddings generated by an unwatermarked LLM and human-written text. The choice of this hyperparameter is critical for the performance of SimMark.
@@
Furthermore, the choice of interval affects the robustness of SimMark against paraphrasing attacks. When an attacker paraphrases sentences, the similarities (or distances) may change and fall outside the interval. A larger interval provides greater robustness, as the watermark is less likely to be disrupted by paraphrasing. Therefore, the selection of the interval involves balancing several factors: it must not be too narrow to impede sentence generation while maintaining a low FP rate and adequate robustness against paraphrasing.
Finding an optimal interval, or “sweet spot,” depends on the distribution of similarities between consecutive sentences, which varies across different embedding models. However, identifying such sweet spots is feasible when the user analyzes the similarity distributions of both LLM-generated and human-written text (See Appendix E for an example of finding the sweet spot).

### v2

# SimMark: A Robust Sentence-Level Similarity-Based Watermarking Algorithm for Large Language Models

@@ Abstract
The widespread adoption of large language models (LLMs) necessitates reliable methods to detect LLM-generated text. We introduce SimMark, a robust sentence-level watermarking algorithm that makes LLMs’ outputs traceable without requiring access to model internals, making it compatible with both open and API-based LLMs. By leveraging the similarity of semantic sentence embeddings combined with rejection sampling to embed detectable statistical patterns imperceptible to humans, and employing a soft counting mechanism, SimMark achieves robustness against paraphrasing attacks. Experimental results demonstrate that SimMark sets a new benchmark for robust watermarking of LLM-generated content, surpassing prior sentence-level watermarking techniques in robustness, sampling efficiency, and applicability across diverse domains, all while maintaining the text quality and fluency. [^1] The source code of our algorithm is available here.
EMNLP’25
SimMark: A Robust Sentence-Level Similarity-Based
Watermarking Algorithm for Large Language Models
Amirhossein Dabiriaghdam  and Lele Wang Department of ECE, University of British Columbia, Vancouver, BC, Canada {amirhossein, lelewang}@ece.ubc.ca
@@ 1 Introduction
The advent of deep generative models has made it increasingly important to determine whether a given text, image, or video was produced by artificial intelligence (AI), and recently, researchers across various domains have begun tackling this challenge Aaronson and Kirchner (2022); Fernandez et al. (2023); Teymoorianfard et al. (2025). In particular, LLMs such as GPT-4o Hurst et al. (2024), can now generate human-like text at scale and low cost, enabling powerful applications across numerous industries and areas such as health care and law Singhal et al. (2023); Wu et al. (2025); Torabi et al. (2025); Yao et al. (2024); Taranukhin et al. (2024).
This capability, however, introduces serious risks, including academic plagiarism, disinformation campaigns, and public opinion manipulation. For instance, the use of AI-generated content in news articles has raised concerns about transparency, accountability, and the spread of false information Futurism (2023). Moreover, reliably detecting LLM-generated content is crucial for enforcing copyright protections and ensuring accountability Weidinger et al. (2021).
Figure 1: A high-level overview of SimMark detection algorithm. The input text is divided into individual sentences $X_{1}$ to $X_{N}$, which are embedded using a semantic embedding model. The similarity between consecutive sentence embeddings is computed. Sentences with similarities within a predefined interval $[a,b]$ are considered valid, while those outside are invalid. A statistical test is performed using the count of valid sentences to determine whether the text is watermarked.
@@
One promising approach is the use of imperceptible statistical signatures, or watermarks, embedded within a text. Watermarking imperceptibly alters text such that it remains natural to human readers but enables subsequent detection of its origin Atallah et al. (2001). Effective watermarking must balance the preservation of the text quality with robustness against adversarial paraphrasing, where attackers modify the text to evade detection Krishna et al. (2024). Additionally, watermarks must be resistant to spoofing attacks, wherein adversaries craft non-machine-generated text (often malicious) to falsely trigger detectors Sadasivan et al. (2023).
In this paper, we introduce SimMark, a robust sentence-level watermarking algorithm for LLMs based on sentence embedding similarity. SimMark treats LLMs as black boxes that can be prompted to generate sentences given a context. This approach makes SimMark compatible with a wide range of models, including open-weight LLMs and closed-source proprietary models accessible only via APIs, as it does not require fine-tuning or access to the models’ internal logits. Access to logits is often restricted by API providers due to their potential use in distilling LLMs and leaking proprietary information Finlayson et al. (2024).
SimMark leverages embeddings from semantic text embedding models to capture semantic relationships between sentences and embeds detectable statistical patterns on sentence similarity through rejection sampling. Specifically, rejection sampling involves querying the LLM multiple times until the similarity between the embeddings of consecutive sentences falls within a predefined interval. During detection, these patterns are analyzed using a statistical test to differentiate between human-written and LLM-generated text, as illustrated in Figure 1.
@@
  • We introduce a novel sentence-level watermarking method that achieves state-of-the-art detection performance while maintaining low false positive rates for human-written text.
@@
  • Compared to existing methods, SimMark provides a more practical solution that operates without access to LLM logits, offering high-quality watermark injection and detection.
The remainder of this paper is organized as follows: Section 2 reviews the background and related work on LLM watermarking techniques. Section 3 outlines our methodology. Section 4 describes our experimental setup and presents comparative results, while Section 5 concludes the paper.
@@ 2.1 Autoregressive Decoding of LLMs
An LLM operates over a vocabulary $V$, a set of words or subwords termed as tokens. Let $f:V\rightarrow V$ be an LLM that takes a sequence of tokens $T_{i}=\{t_{1},t_{2},\ldots,t_{i}\}$ as input and generates the next token $t_{i+1}$ as its output. To generate $t_{i+1}$, the LLM samples it from the conditional probability distribution $P(t_{i+1}|T_{i})$ over the vocabulary $V$. After generating $t_{i+1}$, the updated sequence $T_{i+1}=T_{i}\cup\{t_{i+1}\}$ is fed back into the model, and the process is repeated iteratively to generate the subsequent tokens. This process of generating one token at a time, given the previously generated tokens, is known as autoregressive decoding.
@@ 2.2 Token-Level Watermarking
Token-level watermarking methods embed a statistical signal in the text by manipulating the token sampling process Aaronson and Kirchner (2022); Kirchenbauer et al. (2023); Fu et al. (2024). These methods typically alter the probability distribution over $V$, subtly biasing the selection of certain tokens to form detectable patterns.
KGW introduced by Kirchenbauer et al. (2023), groups $V$ into green and red subsets pseudo-randomly seeded on the previous token before generating each new token. A predefined constant $\delta>0$ is added to the logits of each token in the green list, increasing their likelihood of being selected during the sampling step. At detection, a $z$-test is applied to the number of tokens from the green list in the text to determine whether the text contains a watermark. This test compares the observed proportion of green tokens to the expected proportion under the null hypothesis of no watermark, providing a statistical measure to detect even subtle biases introduced by the watermark.
Detection of such watermarks involves analyzing tokens for statistical signatures that deviate from typical human text. However, token-level watermarks can still be vulnerable to paraphrasing, as rephrasing may disrupt the green and red token lists without altering the overall semantic Krishna et al. (2024). Moreover, since these methods modify the logits, they directly impact the conditional probability distribution over $V$, potentially degrading the quality of the generated text Fu et al. (2024).
Due to space constraints, additional related work—including token-level methods such as UNIGRAM-WATERMARK (UW) Zhao et al. (2023) and the Semantic Invariant Robust (SIR) watermark Liu et al. (2023), as well as post-hoc watermarking techniques—is deferred to Appendix A.
@@ 2.3 Sentence-Level Watermarking
One approach to mitigate the previously mentioned problems is to inject the watermark signal at the sentence level, making it less vulnerable to adversarial modifications Topkara et al. (2006). Consider a similar notation for sentence generation using an autoregressive LLM that takes a sequence of sentences $M_{i}=\{X_{1},X_{2},\ldots,X_{i}\}$ and generates the next sentence $X_{i+1}$. The updated sequence of sentences $M_{i+1}=M_{i}\cup\{X_{i+1}\}$ is then used to generate subsequent sentences iteratively.
@@
SemStamp by Hou et al. (2024a) employs Locality-Sensitive Hashing (LSH) Indyk and Motwani (1998) to pseudo-randomly partition the semantic space of an embedding model into a set of valid and blocked regions, analogous to the green and red subsets in KGW. During rejection sampling, if the embedding of a newly generated sentence lies within the valid regions (determined based on the LSH signature of the previous sentence), the sentence is accepted. Otherwise, a new sentence is generated until a valid sentence is produced or the retry limit is reached. Similar to KGW, a $z$-test is applied to the number of valid sentences to determine whether the text contains a watermark.
@@
To improve robustness against paraphrasing, SemStamp used a contrastive learning approach Hadsell et al. (2006), fine-tuning an embedding model such that the embeddings of paraphrased sentences remain as close as possible to the original sentences. This was achieved by minimizing the distance between paraphrased and original embeddings while ensuring unrelated sentences remained distinct. They also introduce a margin constraint in the rejection sampling process to reject sentences whose embeddings lie near the region boundaries.
@@
$k$-SemStamp Hou et al. (2024b) builds upon SemStamp and aims to enhance robustness by partitioning the semantic space using $k$-means clustering Lloyd (1982) instead of random partitioning. They claim that in this way, sentences with similar semantics are more likely to fall within the same partition, unlike random partitioning, which may place semantically similar sentences into different partitions, reducing robustness; however, $k$-SemStamp assumes that the LLM generates text within a specific domain to apply $k$-means clustering effectively Hou et al. (2024b), limiting its applicability in real-world, open-domain scenarios. The generation and detection procedures of $k$-SemStamp remain similar to the original SemStamp. In contrast to token-level algorithms, these sentence-level methods do not alter the internals of the LLM, therefore, it is expected that their output to be of higher quality Hou et al. (2024a, b).
@@
Our work, similar to SemStamp and $k$-SemStamp, is a sentence-level algorithm; however, it injects its watermark signature into the semantic similarity of consecutive sentences. It achieves great generalizability across domains by leveraging any off-the-shelf, general-purpose embedding model without fine-tuning. At the same time, it outperforms these state-of-the-art (SOTA) sentence-level watermarking methods in robustness against paraphrasing while preserving text quality.
3 SimMark: A Similarity-Based Watermarking Algorithm
@@
Figure 2: Overview of SimMark. Top: Generation. For each newly generated sentence ($X_{i+1}$), its embedding ($e_{i+1}$) is computed using a semantic text embedding model, optionally applying PCA for dimensionality reduction. The cosine similarity (or Euclidean distance) between $e_{i+1}$ and the embedding of the previous sentence ($e_{i}$), denoted as $s_{i+1}$, is calculated. If $s_{i+1}$ lies within the predefined interval $[a,b]$, the sentence is marked valid and accepted. Otherwise, rejection sampling generates a new candidate sentence until validity is achieved or the iteration limit is reached. Once a sentence is accepted, the process repeats for subsequent sentences. Bottom: Detection (+ Paraphrase attack). Paraphrased versions of watermarked sentences are generated ($Y_{i}$), and their embeddings ($e^{\prime}_{i}$) are computed. The similarity between consecutive sentences in the paraphrased text is evaluated. If paraphrasing causes the similarity ($s^{\prime}_{i+1}$) to fall outside $[a,b]$, it is mismarked as invalid. A soft counting mechanism (via function $c(s_{i+1})$ instead of a regular counting with a step function in $[a,b]$) quantifies partial validity based on proximity to the interval bounds, enabling detection of watermarked text via a soft-$z$-test even under paraphrase attacks. It should be noted that soft counting is always applied, as we cannot assume prior knowledge of paraphrasing.
@@ 3.1 Watermarked Text Generation
Similar to SemStamp and $k$-SemStamp, SimMark utilizes the embedding representations of the sentences. To compute the embeddings, in contrast with Hou et al. (2024a, b) that fine-tuned their embedder model (which could make it biased toward a specific paraphrasing model or domain), we employ Instructor-Large Su et al. (2023), a general-purpose embedding model, without any fine-tuning. The flexibility of our method in using any pretrained embedding model enables our approach to be more easily adaptable to different domains.
Algorithm 1 SimMark Generation Pseudo-Code
0: LLM, embedding model, interval $[a,b]$, maximum iterations $N_{\mathrm{max}}$
1:for each generated sentence $X_{i}$ do
@@
5:    Generate sentence $X_{i+1}$ using the LLM.
@@
8:    Compute similarity $s_{i+1}$:
$s_{i+1}\leftarrow\begin{cases}\tfrac{e_{i}\cdot e_{i+1}}{\|e_{i}\|_{2}\cdot\|e_{i+1}\|_{2}}&\text{ if cosine similarity},\\ \|e_{i}-e_{i+1}\|_{2}&\text{ if Euclidean distance,}\end{cases}$
@@
10:  while $s_{i+1}{\color[rgb]{1,0,0}\notin[a,b]}$ and $n<N_{\mathrm{max}}$
11:  Accept $X_{i+1}$ as valid and continue generating the next sentence. {Here we either reached the $N_{\mathrm{max}}$ or $s_{i+1}{\color[rgb]{0,0.5,0}\in[a,b]}$}
@@
First, we compute the embedding for each sentence [^2] In our experiments, we passed both the sentence and “Represent the sentence for cosine similarity:” or “Represent the sentence for Euclidean distance:” as the instruction to the Instructor-Large model.. Then, we calculate the cosine similarity (or Euclidean distance) between the embedding of sentence $i+1$ and the embedding of sentence $i$. If the computed value lies within a predefined interval, sentence $i+1$ is considered valid (analogous to the green subset in KGW). Otherwise, we prompt the LLM to generate a new sentence and repeat this procedure until a valid sentence is found or the maximum number of iterations is reached (in this case, we accept the last generated sentence), as shown in Algorithm 1. Optionally, we may apply Principal Component Analysis (PCA) Jolliffe (2002) method to the embeddings to reduce their dimensionality before calculating the similarity [^3] In our experiments, this is the instruction in this case: “Represent the sentence for PCA:”. The reason for applying PCA is provided in Subsection 3.2. An overview of SimMark generation algorithm is depicted in the top part of Figure 2.
The predefined interval is a hyperparameter chosen a priori based on the distribution of similarities between consecutive sentences’ embeddings generated by an unwatermarked LLM and human-written text. The choice of this hyperparameter is critical for the performance of SimMark.
@@
Furthermore, the choice of interval affects the robustness of SimMark against paraphrasing attacks. When an attacker paraphrases sentences, the similarities may change and fall outside the interval. Consequently, a larger interval provides greater robustness, as the watermark is less likely to be disrupted by paraphrasing. Therefore, the selection of the interval involves balancing several factors: it must not be too narrow to impede sentence generation while maintaining a low FP rate and adequate robustness against paraphrasing.
Finding a “sweet spot,” for the interval depends on the distribution of similarities between consecutive sentences, which can vary across models. However, identifying such sweet spots is feasible when we analyze the similarity distributions of both human-authored and LLM-generated text (see Appendix K for an example of finding a sweet spot).